OpenStudy (anonymous):

The tip of a 15-inch wiper blade wipes a path that is 35 inches long. What is the angle of rotation of the blade in radians to the nearest tenth?

OpenStudy (anonymous):

How would I do this? Sorry to bother you again - @hartnn

OpenStudy (anonymous):

I just need a formula

hartnn (hartnn):

here s = length of arc = 35 r = 15 use \(\Large s = r \theta \)

hartnn (hartnn):

and you'll get theta in radians only :)

OpenStudy (anonymous):

so s/r = 2.33

hartnn (hartnn):

35 = 15 theta theta = 35/15 = 7/3 radians or 2.33 radians

hartnn (hartnn):

thats the angle of rotation :)